# Edge Graph Intelligence: Reciprocally Empowering Edge Networks with Graph Intelligence

### v1

@@ VI-D GI-Assisted Federated Learning
In certain FL scenarios, clients exhibit relational dependencies, a structural representation of which can be established through a graph. For instance, traffic sensors in physical proximity often result appears in similar traffic patterns. In this case, a cross-client graph can be established by identifying them as respective vertices. The presence of this cross-client graph prompts the utilization of GI algorithms to enhance the FL training process, resulting in the paradigm of GI-assisted FL (GFL), as shown in Fig. 13. As we have discussed in Sec. V-A, GFL and FGL are different mechanisms, where the former applies GI techniques to optimize FL on arbitrary ML models (which do not necessarily to be GI models) and the latter applies FL to jointly train GI models. In other words, GI models are optimization tools in GFL, whereas they are the targeted models to be trained in FGL. Due to such distinct functionality, we categorize GFL into the GI-based optimization for edge networks and discuss it in this subsection, while FGL will be introduced in the context of edge computation of GI models in Sec. V-A.
@@
GFL works upon the premise that clients closely connected in the graph likely share similar data distributions. Based on that, GNNs can utilize the graph structure inherent in the FL system to tackle the non-IID data hetergeneity among clients. With respect to the utilization of centralized servers, we discuss GFL in the centralized setting and the decentralized setting, respectively.
VI-D1 Centralized Federated Learning
@@
For the case of server-side GI execution, a GI model undergoes training on the server using graph relationships between clients, operating under the assumption that proximate clients exhibit similarities in their local models or feature embeddings. Initially, the server initiates the collection of parameters from clients, following a standard FL protocol. These uploaded parameters collected from client models are construed as node features within the context of the cross-client graph, using which the server trains a GI model to facilitate the aggregation process in FL. Note that the specification of the cross-client graph may either be predefined or dynamically extracted through the application of a self-attention module during the training phase [283]. The resultant updated parameters are then transmitted back to the respective clients. In this process, researchers have proposed different routines to tackle different issues in training the coordinated GI model. For instance, some researchers focus on the convergence speed of GI models under non-IID circumstances, designing bi-level optimization to train both the client models and the server-side GI model simultaneously, e.g., using unsupervised contrastive learning [284] or regularization-aided supervised learning methods [285]. Others ponder the specific functionality of the GI models and develop training strategies in various aspects [286, 287]. PDGNet [286] designs a primal-dual iterative approach to train a GNN model for minimizing the transmission error probability across clients, in order to find the optimal power allocation policy among edge networks. CNNFGNN [288] performs alternating optimization that updates clients’ model weights and the GI model weights by freezing either of them by turns.
For the case of client-side GI execution, each client is assumed to possess a global cross-client graph and trains not only its local model but a GI model to obtain global knowledge from other clients. Building different graphs within the clients can be employed to address various problem-solving objectives like data heterogeneity and model heterogeneity. FedCG [289] establishes a fully connected graph by leveraging the similarity among clients’ model weights or pattern features. Each client undertakes the training of a GNN using the cross-client graph and derives a global embedding. Subsequently, the local model embedding and the global embedding are amalgamated through the application of a trainable weight, which enhances the capacity of the model to incorporate both local and global information, addressing data heterogeneity. HAFL-GHN [290] treats neural architectures within individual clients as graphs, where model layers are vertices and their connections are links. On these graphs, HAFL-GHN develops a graph hyper-network to minimize empirical risk across all clients, so as to facilitate indirect aggregation of heterogeneous model weights across clients.
@@
VI-D2 Decentralized Federated Learning
@@
Weight summation methods update clients’ local models through aggregating their neighbors’ local model parameters following the global cross-client graph topology. Under this principle, DSGT [291] employs decentralized stochastic gradient tracking to achieve faster convergence, and PSO-GFML [292] selectively exchanging local model parameters with the servers to improves communication efficiency. The weights of the cross-client graph’s links is determined based on the similarity between unlabeled graph embeddings [293] or hidden parameters [294].
@@
Graph regularization integrates graph Laplacian regularization into the objective function to transform model parameters from neighboring clients [295] and is particularly beneficial to multi-task GFL. In dFedU [296], a pre-defined fully connected cross-client graph is given ahead where each client is assigned a single task. When receiving local updated models from neighboring clients, each client conducts model updating with graph regularization. He et al. [297] alternatively posit that each client runs multiple tasks and initialize a cross-client task relationship graph with task classifier parameters, where decentralized periodic averaging SGD is employed to optimize the objective function. Wang et al. [298] proposes to learn each client’s local loss function through a decentralized alternating direction method of multipliers algorithm with statistical guarantees.
VII Edge Graph Intelligence Ecosystems
The closed loop of edge networks and graph intelligence have profitably retrofitted each other, inflating an abundant ecosystem to sustain EGI. Fig. 14 shows the landscape of the EGI ecosystem from levels of hardware, software, benchmark, and applications. Given the applications introduced in Sec. IV, we discuss hardware, software (i.e., development frameworks), and benchmarks in what follows.
VII-A Edge Hardware for EGI
@@
Edge networks cover a wide spectrum of edge platforms from embedded and mobile devices to vehicles and edge servers, as discussed in Sec. III-B. At the core of these edge facilities, different processors deliver different capabilities and require different optimizations for GI model computation. In general, they can be discussed in three types, mobile CPUs and GPUs, Field Programmable Gate Array (FPGA), and Domain-Specific Accelerators (DSAs).
VII-A1 CPU and GPU
@@
A majority of edge facilities, especially wearable and mobile devices, are equipped only with conventional CPUs and GPUs such as Qualcomm SnapDragon series and Huawei Kirin series. Given the intensive workload of GI-driven applications, efficient execution of them requires meticulously exploiting the capability of low-power processors. For better understanding the characteristics of GI models on conventional platforms, Huang et al. [299] examine in detail representative GI systems and reveal five major gaps in optimizing GI computation performance, i.e, poor data locality, severe workload imbalance, redundant memory access, large memory footprint, and inefficiency on variant feature lengths. Based on these observations, Huang et al. [299] propose a set of tailored optimizations to bridge the gaps and achieve 1.37$\times$-15.5$\times$ performance improvement on various GI models. Towards the same objective, Zhang et al. [300] study the GI execution performance on commercial platforms from the perspective of computational graph and summarize three issues that existing systems suffer, namely redundant neural operator computation, inconsistent thread mapping, and excessive intermediate data. To tackle these challenges, Zhang et al. [300] introduce three corresponding designs, i.e., propagation-postponed operator reorganization, unified thread mapping for fusion, and intermediate data recomputation, achieving notable speedup in lower memory IO and footprint.
@@
These empirical studies motivate a number of literature [301, 302, 303, 304, 305] in efficiently leveraging the potential of conventional CPUs and GPUs. For instance, Adiletta et al. [306] study a detailed characterization of GI models on Intel’s Programmable Integrated Unified Memory Architecture (PIUMA) for scalable training. Graphite [302] addresses GI model training on CPUs by eliminating the memory limitation through software-hardware co-design including layer fusion, feature compression, and a vertex processing reordering algorithm for improving temporal locality, and implements a high-performance GI computation system on Intel CPUs. Some works explore optimizing matrix computation in GI tasks from the kernel level. FeatGraph [307] provides efficient implementations of Sampled Dense-Dense Matrix Multiplication (SDDMM) and Sparse-dense Matrix Multiplication (SpMM) for both CPUs and GPUs, and FusedMM [301] further develops a general solution for different GI programming backends. There are also many comprehensive systems developed to tackle the various aspects of GI training and inference for distributed training on CPU or GPU clusters (e.g., Dorylus [304], FlexGraph [305]). However, most of them rely on the computation infrastructure at datacenter levels and are thus out of the scope of this survey. Interested readers may refer to recent reviews [308, 309] on large-scale distributed GI systems.
@@
VII-A2 Field-Programmable Gate Array
Field-Programmable Gate Array is a type of integrated circuit with a differentiated ability of being (re)programmed to desired functionality requirements after manufacturing. This advanced ability allows FPGA to be used in applications with requirements on flexibility, speed, and parallelism, and thus are embedded and employed in edge platforms for GI workload processing. HP-GNN [310] generates high-throughput GNN training implementations on a given CPU-FPGA platform by designing an automatic algorithm that maps GI training algorithms, GI models, and hardware specifics to the targeted CPU-FPGA platform. GenGNN [311] develops a generic GI model acceleration framework using High-Level Synthesis (HLS) techniques, aiming to render GI inference in low latency and support various GI models in flexible extensibility. It embraces an optimized message-passing structure to accommodate the majority of popular GI models and provides a rich library of model-specific components. GraphAGILE [312] designs a hardware module named adaptive computation kernel for efficient execution of GI matrix operations such as sparse-dense matrix multiplication and sampled dense-dense matrix multiplication, and proposes a set of compiler optimizations to reduce inference latency. Graph-OPU [313] proposes an FPGA-based overlay processor for accelerating GI execution, which introduces microarchitecture for fully-pipelined GI processing and customizes the instruction sets for various GI model adoptions.
@@ VII-A2 Field-Programmable Gate Array
VII-A3 Domain-Specific Accelerator
@@
Domain-Specific Accelerators (DSAs) are important computing components and becoming more pervasive in many edge facilities such as dedicated edge servers and some mobile devices [317, 318, 319, 320]. By optimizing specifically for a narrow range of computations, DSAs are tailored to meet the needs of targeted algorithms in dedicated domains and enable orders of magnitude improvements in performance or cost compared to general-purpose processors. Motivated by the achieved success of DSAs in CNNs, the community has made plentiful efforts in developing DSAs for GI workload. Towards that, EnGN [321] presents a unified architecture inspired by CNN accelerators, where GI computations are viewed in a matrix perspective and matrix multiplication optimizations are applied for improved runtime performance. Auten et al. [322] propose a modular DSA architecture for convolutional GNNs, where each tile as a basic unit of it composes of an aggregation module, a DNN accelerator module, a DNN queue, and a graph processing element.
@@
Nonetheless, GNNs are semantically different from traditional DNNs (c.f. Sec. II), bringing new computing characteristics desired for tailored designs. To better understand such computing characteristics, Yan et al. [323] and Lin et al. [324] respectively conduct detailed profiling of general GNN workload and distributed GNN training workload on GPUs, providing useful insight on GNN-oriented DSA design. Based on that, HyGCN [325] introduces a hybrid architecture to tame the alternating phases of GI computation, which is composed of separate dedicated engines for the aggregate and update functions as well as a control scheme to pipeline the execution of both functions. GRIP [326] leverages the programming abstraction of GReTA [327] to develop a general DSA for GNNs, which organizes a GI model inference into four steps, i.e., gather, reduce, transform, and activate. Regarding this abstraction, GRIP specializes the processing units for links and vertices separately and designs a control module to coordinate data movement between units and buffers. AWB-GCN [328] raises an aggressive adaptation to the structural sparsity of GI models under the motivation that for power-law graphs, GI computation can be polarly dense or sparse and suffers from workload imbalance. To address this imbalance, AWB-GCN [328] develops a custom matrix multiplication engine with autotuning workload balancing techniques, and GNNIE advocates a flexible MAC architecture that splits features into blocks with load (re)distribution and graph-specific caching to bypass the high costs of random DRAM.
VII-B Developmemt Frameworks
@@
GI models’ unique computing workflow introduces different programming patterns from traditional DL models (e.g., CNNs and RNNs), which requires dedicated frameworks for implementation. Towards that, the community has developed various programming frameworks compatible with edge computing platforms. Among them, PyTorch Geometric (PyG) [329] and Deep Graph Library (DGL) [330] are the two most popular ones and are widely used in various applications. PyG is a geometric deep learning extension library built upon PyTorch that provides flexible interfaces for programming GI models on both CPUs and GPUs. It adopts a message-passing programming abstraction for building GI models, which is programmer-friendly to express various GI variants. DGL is another general framework specialized for deep learning models on graphs. It also leverages the message-passing primitives in a user-configurable way and introduces a set of parallelization strategies for high-performance GNN execution. Both PyG and DGL affiliate model zoos for a quick start of implementation. Besides PyG and DGL, the industry has also developed many frameworks upon their customized demand, such as AGL [331], BGL [332], TF-GNN [333], and Angel-Graph, etc. Nonetheless, most of them are dedicated to cloud-level resources without consideration of the deployment on edge networks and thus require tailored adjustments when building edge GI services.
@@
Intelligent Robotics & Vehicles US-101
@@
VII-C Open Datasets
To assess how well edge GI systems perform and learn their behaviors upon deployment, versatile datasets are collected from diverse edge scenarios, serving as performance benchmarks. Table III lists some representative datasets in several application domains discussed in Sec. IV.
The first dataset series lies on smart cities, including data collected from intelligent transportation systems, meteorological stations, power grids, and governments. Their graphs are mostly derived from the geographical network relations in cities, e.g., road nets, power networks, and regional adjacencies. The second domain is intelligent robotics and vehicles, taking datasets from various robot manipulation scenarios. In particular, Gym from Open AI provides a high-fidelity simulation environment for robot learning, which can be well utilized for collecting necessary data needed by GI model training and testing. The third category is for human sensing applications, including human action recognition, facial affective analysis, and sleep quality prediction. Specifically, for the sleep quality dataset, researchers construct sleep stage graphs with electronic channel records and employ GI models for prediction. The fourth domain is location-based social recommendation, where social networks are naturally graphs and the location information of users puts its analysis into the context of edge networks. The last group focuses on mobile vision, where the datasets are also widely utilized in many computer vision models. Readers are encouraged to exact their interested datasets from the links to learn more about their sources and applications.
VIII Open Challenges and Future Research Opportunities

### v2

@@ VII-D GI-Assisted Federated Learning
In certain FL scenarios, clients exhibit relational dependencies, a structural representation of which can be established through a graph. For instance, traffic sensors in physical proximity often result appear in similar traffic patterns. In this case, a cross-client graph can be established by identifying them as respective vertices. The presence of this cross-client graph prompts the utilization of GI algorithms to enhance the FL training process, resulting in the paradigm of GI-assisted FL (GFL), as shown in Fig. 16. As we have discussed in Sec. VI-A, GFL and FGL are different mechanisms, where the former applies GI techniques to optimize FL on arbitrary ML models (which do not necessarily be GI models) and the latter applies FL to jointly train GI models. In other words, GI models are optimization tools in GFL, whereas they are the targeted models to be trained in FGL. Due to such distinct functionality, we categorize GFL into the GI-based optimization for edge networks and discuss it in this subsection, while FGL will be introduced in the context of edge computation of GI models in Sec. VI-A.
@@
GFL works upon the premise that clients closely connected in the graph likely share similar data distributions. Based on that, GNNs can utilize the graph structure inherent in the FL system to tackle the non-IID data heterogeneity among clients. With respect to the utilization of centralized servers, we discuss GFL in the centralized setting and the decentralized setting, respectively.
VII-D1 Centralized Federated Learning
@@
For the case of server-side GI execution, a GI model undergoes training on the server using graph relationships between clients, operating under the assumption that proximate clients exhibit similarities in their local models or feature embeddings. Initially, the server initiates the collection of parameters from clients, following a standard FL protocol. These uploaded parameters collected from client models are construed as node features within the context of the cross-client graph, using which the server trains a GI model to facilitate the aggregation process in FL. Note that the specification of the cross-client graph may either be predefined or dynamically extracted through the application of a self-attention module during the training phase [272]. The resultant updated parameters are then transmitted back to the respective clients.
In this process, researchers have proposed different routines to tackle different issues in training the coordinated GI model. For instance, some researchers focus on the convergence speed of GI models under non-IID circumstances, designing bi-level optimization to train both the client models and the server-side GI model simultaneously, e.g., using unsupervised contrastive learning [273] or regularization-aided supervised learning methods [274]. Others ponder the specific functionality of the GI models and develop training strategies in various aspects [275, 276].
@@
For the case of client-side GI execution, each client is assumed to possess a global cross-client graph and trains not only its local model but a GI model to obtain global knowledge from other clients. Building different graphs within the clients can be employed to address various problem-solving objectives like data heterogeneity and model heterogeneity. For instance, FedCG [277] establishes a fully connected graph by leveraging the similarity among clients’ model weights or pattern features. Each client undertakes the training of a GNN using the cross-client graph and derives a global embedding. Subsequently, the local model embedding and the global embedding are amalgamated by applying a trainable weight, which enhances the capacity of the model to incorporate both local and global information, addressing data heterogeneity.
VII-D2 Decentralized Federated Learning
@@
Weight summation methods update clients’ local models by aggregating their neighbors’ local model parameters following the global cross-client graph topology. Under this principle, DSGT [278] employs decentralized stochastic gradient tracking to achieve faster convergence, and PSO-GFML [279] selectively exchanging local model parameters with the servers to improve communication efficiency. The weights of the cross-client graph’s links are determined based on the similarity between unlabeled graph embeddings [280] or hidden parameters [281].
@@
Graph regularization integrates graph Laplacian regularization into the objective function to transform model parameters from neighboring clients [282] and is particularly beneficial to multi-task GFL. In dFedU [283], a pre-defined fully connected cross-client graph is given ahead where each client is assigned a single task. When receiving local updated models from neighboring clients, each client conducts model updating with graph regularization. He et al. [284] alternatively posit that each client runs multiple tasks and initializes a cross-client task relationship graph with task classifier parameters, where decentralized periodic averaging SGD is employed to optimize the objective function.
Lessons Learned (Sec. VII) GI can serve as an optimization tool for enhancing different functionalities of edge networks, including network resource management, graph-based decision-making, security and anomaly detection, and GFL. To well harness GI in these cases, there are dual knobs: on the one hand, the system needs to concisely identify the optimization constraints and objectives from graph perspectives, where one or more graph(s) should be built to represent the structure of the targeted problem; on the other hand, the graph may associate appropriate (encoded) properties that supply sufficient information and convert optimization objectives into learning objectives, such that GI models can use these graph semantics to induce expected output.
VIII Edge Graph Intelligence Ecosystems
After reviewing the interaction between GI and edge networks in Sec. VI and Sec. VII, we discuss the EGI ecosystem inflated from this loop. Fig. 17 shows the landscape of the EGI ecosystem from levels of hardware, software, benchmark, and applications.
VIII-A Edge Hardware for EGI
@@
Edge networks cover a wide spectrum of edge platforms from embedded and mobile devices to vehicles and edge servers, as discussed in Sec. III-B2. At the core of these edge facilities, different processors deliver different capabilities and require different optimizations for GI model computation. In general, they can be discussed in three types, mobile CPUs and GPUs, Field Programmable Gate Array (FPGA), and Domain-Specific Accelerators (DSAs).
VIII-A1 CPU and GPU
@@
A majority of edge facilities, especially wearable and mobile devices, are equipped only with conventional CPUs and GPUs such as Qualcomm SnapDragon series and Huawei Kirin series. Given the intensive workload of GI-driven applications, efficient execution of them requires meticulously exploiting the capability of low-power processors. To better understand the characteristics of GI models on conventional platforms, Huang et al. [285] examine in detail representative GI systems and reveal five major gaps in optimizing GI computation performance, i.e, poor data locality, severe workload imbalance, redundant memory access, large memory footprint, and inefficiency on variant feature lengths. Based on these observations, Huang et al. [285] propose a set of tailored optimizations to bridge the gaps and achieve 1.37$\times$-15.5$\times$ performance improvement on various GI models. Towards the same objective, Zhang et al. [286] study the GI execution performance on commercial platforms from the perspective of computational graph and summarize three issues that existing systems suffer, namely redundant neural operator computation, inconsistent thread mapping, and excessive intermediate data. To tackle these challenges, Zhang et al. [286] introduce three corresponding designs, i.e., propagation-postponed operator reorganization, unified thread mapping for fusion, and intermediate data recomputation, achieving notable speedup in lower memory IO and footprint.
@@
These empirical studies motivate literature [287, 288, 289, 290, 291] in efficiently leveraging the potential of conventional CPUs and GPUs. For instance, Adiletta et al. [292] study a detailed characterization of GI models on Intel’s Programmable Integrated Unified Memory Architecture (PIUMA) for scalable training. Graphite [288] addresses GI model training on CPUs by eliminating the memory limitation through software-hardware co-design including layer fusion, feature compression, and a vertex processing reordering algorithm for improving temporal locality, and implements a high-performance GI computation system on Intel CPUs. Some works explore optimizing matrix computation in GI tasks from the kernel level. FeatGraph [293] provides efficient implementations of Sampled Dense-Dense Matrix Multiplication (SDDMM) and Sparse-dense Matrix Multiplication (SpMM) for both CPUs and GPUs, and FusedMM [287] further develops a general solution for different GI programming backends. There are also many comprehensive systems developed to tackle the various aspects of GI training and inference for distributed training on CPU or GPU clusters (e.g., Dorylus [290], FlexGraph [291]). However, most of them rely on the computation infrastructure at datacenter levels and are thus out of the scope of this survey. Interested readers may refer to recent reviews [294, 295] on large-scale distributed GI systems.
@@
VIII-A2 Field-Programmable Gate Array
FPGA is a type of integrated circuit with a differentiated ability to be (re)programmed to desired functionality requirements after manufacturing. This advanced ability allows FPGA to be used in applications with requirements on flexibility, speed, and parallelism, and thus are embedded and employed in edge platforms for GI workload processing. HP-GNN [296] generates high-throughput GNN training implementations on a given CPU-FPGA platform by designing an automatic algorithm that maps GI training algorithms, GI models, and hardware specifics to the targeted CPU-FPGA platform. GenGNN [297] develops a generic GI model acceleration framework using High-Level Synthesis (HLS) techniques, aiming to render GI inference in low latency and support various GI models in flexible extensibility. It embraces an optimized message-passing structure to accommodate the majority of popular GI models and provides a rich library of model-specific components. GraphAGILE [298] designs a hardware module named adaptive computation kernel for efficient execution of GI matrix operations such as sparse-dense matrix multiplication and sampled dense-dense matrix multiplication, and proposes a set of compiler optimizations to reduce inference latency. Graph-OPU [299] proposes an FPGA-based overlay processor for accelerating GI execution, which introduces microarchitecture for fully-pipelined GI processing and customizes the instruction sets for various GI model adoptions.
@@ VIII-A2 Field-Programmable Gate Array
VIII-A3 Domain-Specific Accelerator
@@
Domain-Specific Accelerators (DSAs) are important computing components and becoming more pervasive in many edge facilities such as dedicated edge servers and some mobile devices [303, 304, 305, 306]. By optimizing specifically for a narrow range of computations, DSAs are tailored to meet the needs of targeted algorithms in dedicated domains and enable orders of magnitude improvements in performance or cost compared to general-purpose processors. Motivated by the achieved success of DSAs in CNNs, the community has made plentiful efforts in developing DSAs for GI workload. Towards that, EnGN [307] presents a unified architecture inspired by CNN accelerators, where GI computations are viewed in a matrix perspective and matrix multiplication optimizations are applied for improved runtime performance. Auten et al. [308] propose a modular DSA architecture for convolutional GNNs, where each tile as a basic unit is composed of an aggregation module, a DNN accelerator module, a DNN queue, and a graph processing element.
@@
Nonetheless, GNNs are semantically different from traditional DNNs (c.f. Sec. III-A), bringing new computing characteristics desired for tailored designs. To better understand such computing characteristics, Yan et al. [309] and Lin et al. [310] respectively conduct detailed profiling of general GNN workload and distributed GNN training workload on GPUs, providing useful insight on GNN-oriented DSA design. Based on that, HyGCN [311] introduces a hybrid architecture to tame the alternating phases of GI computation, which is composed of separate dedicated engines for the aggregate and update functions as well as a control scheme to pipeline the execution of both functions. GRIP [312] leverages the programming abstraction of GReTA [313] to develop a general DSA for GNNs, which organizes a GI model inference into four steps, i.e., gather, reduce, transform, and activate. Regarding this abstraction, GRIP specializes in the processing units for links and vertices separately and designs a control module to coordinate data movement between units and buffers. AWB-GCN [314] raises an aggressive adaptation to the structural sparsity of GI models under the motivation that for power-law graphs, GI computation can be polarly dense or sparse and suffers from workload imbalance. To address this imbalance, AWB-GCN [314] develops a custom matrix multiplication engine with autotuning workload balancing techniques, and GNNIE [315] advocates a flexible MAC architecture that splits features into blocks with load (re)distribution and graph-specific caching to bypass the high costs of random DRAM.
VIII-B Developmemt Frameworks
@@
GI models’ unique computing workflow introduces different programming patterns from traditional DL models (e.g., CNNs and RNNs), which requires dedicated frameworks for implementation. Towards that, the community has developed various programming frameworks compatible with edge computing platforms. Among them, PyTorch Geometric (PyG) [316] and Deep Graph Library (DGL) [317] are the two most popular ones and are widely used in various applications. PyG is a geometric deep learning extension library built upon PyTorch that provides flexible interfaces for programming GI models on both CPUs and GPUs. It adopts a message-passing programming abstraction for building GI models, which is programmer-friendly to express various GI variants. DGL is another general framework specialized for deep learning models on graphs. It also leverages the message-passing primitives in a user-configurable way and introduces a set of parallelization strategies for high-performance GNN execution. Both PyG and DGL can be used in edge devices and edge servers as long as they enable the PyTorch library in their execution environments. In addition, DGL also supports TensorFlow and MXNet, and edge platforms that run these two frameworks also allow model execution with DGL. Besides PyG and DGL, the industry has also developed many frameworks upon their customized demand, such as AGL [318], BGL [319], TF-GNN [320], and Angel-Graph, etc. Nonetheless, most of them are dedicated to cloud-level resources without consideration of the deployment on edge networks and thus require tailored adjustments when building edge GI services.
@@
Intelligent Robots & Vehicles US-101
@@
nuScenes
Vertex: Vehicles or pedestrians
Link: Entity relationship
A large-scale, multimodal dataset for autonomous driving, containing annotated sensor data from LiDAR, radar, cameras, and GPS across diverse driving scenarios in urban environments. [332]
CORSMAL
Vertex: Entity
Link: Entity relationship
A multimodal dataset focused on robotic perception for manipulating containers, including audio, visual, and depth data to estimate container properties and contents in real-time interactions. [333]
RoboNet
Vertex: Entity
Link: Entity relationship
Collection of robotic manipulation data across diverse objects, environments, and robotypes, designed to enable learning generalizable manipulation policies. [334]
@@
VIII-C Open Datasets
To assess how well edge GI systems perform and learn their behaviors upon deployment, versatile datasets are collected from diverse edge scenarios, serving as performance benchmarks. Table III lists some representative datasets in several application domains discussed in Sec. V.
The first dataset series lies on smart cities, including data collected from intelligent transportation systems, meteorological stations, power grids, and governments. Their graphs are mostly derived from the geographical network relations in cities, e.g., road nets, power networks, and regional adjacencies. The second domain is intelligent robotics and vehicles, taking datasets from various robot manipulation scenarios such as UAV swarms and vehicle trajectories The third category is for human sensing applications, including human action recognition, facial affective analysis, and sleep quality prediction. Specifically, for the sleep quality dataset, researchers construct sleep stage graphs with electronic channel records and employ GI models for prediction. The fourth domain is location-based social recommendation, where social networks are naturally graphs and the location information of users puts its analysis into the context of edge networks. The last group focuses on mobile vision, where the datasets are also widely utilized in many computer vision models. Readers are encouraged to exact their interested datasets from the links to learn more about their sources and applications.
Lessons Learned (Sec. VIII) The promising fusion of EGI facilitates the EGI ecosystem in a full-stack manner. Specifically, the ecosystem provides an abundant set of facilities covering hardware, software, and testing benchmarks, acting as a stage for developing and deploying EGI services. Developers are suggested to develop their EGI services agilely with well-established open-sourced toolkits.
IX Open Challenges and Future Research Opportunities